# Around the ‘‘Fundamental Theorem of Algebra’’ (extended version)

### v1

@@ Abstract.
This text is a supplemented and more detailed version of [33].
@@ A Laurent polynomial in $n$ variables is called real if it takes real values on the compact subtorus
Indeed, for compact convex sets $A,B$, the inclusion $A\subseteq B$ is equivalent to the inequality $h_{A}\leq h_{B}$ of their support functions. For $\xi\in{\mathbb{R}}^{n}$, the support function of ${\rm conv}(\Lambda)$ is
@@ 3.3. Newton ellipsoid of group representation
(3) If the group $K$ is simple, then the ellipsoid ${\rm Ell}(\pi)$ is a ball in $\mathfrak{k}^{*}$ for any coadjointly invariant metric in $\mathfrak{k}^{*}$.
@@ 4.1. Exponential sums of one variable
$f(z)=\sum\nolimits_{\lambda\in\Lambda\subset{\mathbb{C}},\,c_{\lambda}\in{\mathbb{C}}}c_{\lambda}\>{\rm e}^{\bar{\lambda}z}.$
The set $\Lambda$ is called the spectrum of $f$, and the convex hull $\Delta={\rm conv}(\Lambda)$ is referred to as the Newton polygon of $f$. If $\Lambda>1$, the zero set of an exponential sum is infinite. For example, the function $f(z)=\>{\rm e}^{2\pi iz}-1$ vanishes precisely on $\mathbb{Z}$}.
The following result may be regarded as an analogue of the Fundamental Theorem of Algebra for exponential sums.
@@
Let $N(f,r)$ denote the number of zeros of an exponential sum $f$ inside the disk $B_{r}$ of radius $r$ centered at the origin. Then
@@
If the convex polygon $\Delta$ degenerates to a line segment, its perimeter tends to twice the length of that segment. Hence, when the Newton polygon of $f$ is a segment, the quantity $l(f)$ in (4.1) equals the length of this segment.
@@
Under the substitution $z\mapsto\>{\rm e}^{z}$, a Laurent polynomial becomes an exponential sum. Therefore, Theorem 4.1 implies the classical Fundamental Theorem of Algebra for polynomials.
Proof of Theorem 4.1. Let $r_{1},\ldots,r_{N}$ be the rays generated by the outward normals to the sides $\Delta_{1},\ldots,\Delta_{N}$ of the Newton polygon $\Delta$ associated with $f$. For each $j$, let $U_{j,R}$ denote the set of points in $\mathbb{C}$} whose distance from $r_{j}$ does not exceed $R$. If $R$ is sufficiently large, then all zeros of $f$ are contained in $\bigcup_{1\leq j\leq N}U_{j,R}$. Denote by $N_{j}(f,r)$ the number of zeros of $f$ lying in $U_{j,R}\cap B_{r}$. Using the argument principle, it is not difficult to obtain
$N_{j}(f,r)=\frac{r}{2\pi}\,l_{j}(f)+o(r),$ (4.2)
where $l_{j}(f)$ is half the length of the side $\Delta_{j}$. Summing over all $j$ yields the leading term of the asymptotic relation (4.1). It remains to show that the $o(r)$ term in (4.2) can, in fact, be replaced by an $O(1)$ term. This follows from the next lemma.
@@
Let $f$ be an exponential sum, and let $A\subset{\mathbb{C}}$ be a compact set. Then there exists a constant $C>0$ such that, for all $z\in{\mathbb{C}}$, the number of zeros of $f$ inside the translated set $z+A$ is bounded above by $C$.
@@
Consider the space of exponential sums with a fixed spectrum $\Lambda$, and let $\mathbb{P}$} denote its projectivization. We identify points of $\mathbb{P}$} with exponential sums defined up to multiplication by a nonzero scalar. Under this identification, the shift action of $\mathbb{C}$} on exponential sums, given by $w\colon g(z)\mapsto g(z+w)$, induces a projective representation
$\rho\colon{\mathbb{C}}\to\mathrm{Aut}({\mathbb{P}})$
of the additive group $\mathbb{C}$} in the projective space $\mathbb{P}$}.
To obtain a contradiction, suppose that such a constant $C$ does not exist. Then there would exist a sequence of points $w_{1},w_{2},\ldots\in{\mathbb{C}}$ such that the number of zeros of the function $f(w_{k}+z)$ inside $A$ tends to infinity as $k\to\infty$. We can choose this sequence so that the corresponding sequence $\rho(w_{k})(f)$ converges in $\mathbb{P}$} to a limit $f_{\infty}$. However, this would imply that $f_{\infty}$ has infinitely many zeros in $A$, which is impossible. ∎
Let $A$ be a rectangle in $\mathbb{C}$} whose one side has length exceeding $2R$ and is perpendicular to the ray $r_{j}$. Let $M(z)$ denote the number of zeros of $f$ in the domain $z+A$. It follows from the argument principle that for any $j$
$\exists C\colon|N_{j}(f,r)-\frac{r}{2\pi}\,l_{j}(f)|\leq C+\sup_{z\in{\mathbb{C}}}M(z).$
Then, applying Lemma 4.1, we obtain that the remainder term in (4.2) admits the estimate $O(1)$.
@@ 4.2.1. Definitions and examples
Assume that the spectra $\Lambda_{1},\ldots,\Lambda_{n}$ are contained in ${\rm Re\>}({{\mathbb{C}}^{n}}^{*})$. Then, for any domain $D\subset{\rm Im\>}({\mathbb{C}}^{n})$,

### v2

# Around the "Fundamental Theorem of Algebra" (extended version)

@@ A Laurent polynomial in $n$ variables is called real if it takes real values on the compact subtorus
Indeed, for centrally symmetric compact convex sets $A,B$, the inclusion $A\subseteq B$ is equivalent to the inequality $h_{A}\leq h_{B}$ of their support functions. For $\xi\in{\mathbb{R}}^{n}$, the support function of ${\rm conv}(\Lambda)$ is
@@ 3.3. Newton ellipsoid of group representation
(3) If the group $K$ is simple, then the ellipsoid ${\rm Ell}(\pi)$ is a ball in $\mathfrak{k}^{*}$ for any coadjointly invariant metric in $\mathfrak{k}^{*}$. (This property of the ellipsoid ${\rm Ell}(\pi)$ is used in calculating the asymptotics of the number of real roots in Theorem 3.3.)
@@ 4.1. Exponential sums of one variable
$f(z)=\sum_{\lambda\in\Lambda}c_{\lambda}\,\>{\rm e}^{\bar{\lambda}z},$
where $\Lambda\subset{\mathbb{C}}$ is a finite set and $c_{\lambda}\in{\mathbb{C}}$. The set $\Lambda$ is called the spectrum of $f$, and its convex hull
$\Delta={\rm conv}(\Lambda)$
is referred to as the Newton polygon of $f$. If $\Lambda>1$, the zero set of an exponential sum is infinite; for example, the function $f(z)=\>{\rm e}^{2\pi iz}-1$ vanishes precisely on $\mathbb{Z}$}.
The following result may be viewed as an analogue of the Fundamental Theorem of Algebra for exponential sums.
@@
Let $N(f,r)$ denote the number of zeros of an exponential sum $f$ in the disk $B_{r}=\{z\in{\mathbb{C}}:\,|z|\leq r\}$. Then
@@
If the polygon $\Delta$ degenerates to a line segment, its perimeter converges to twice the length of that segment. In this case, the quantity $l(f)$ in (4.1) coincides with the length of $\Delta$.
@@
Under the substitution $z\mapsto\>{\rm e}^{z}$, a Laurent polynomial is transformed into an exponential sum. Thus Theorem 4.1 recovers the classical Fundamental Theorem of Algebra.
We begin the proof of Theorem 4.1 with an auxiliary construction. Let $K\subset{\mathbb{C}}$ be a smooth compact curve. Assume first that $K$ does not intersect the zero set of an exponential sum $g$. Let ${\rm Arg}_{g,K}$ denote a branch of the multivalued function $\arg g(z)$ defined along $K$. The differential $d{\rm Arg}_{g,K}$ is independent of the chosen branch. We set
${\rm Arg}_{g}(K)=\int_{K}d{\rm Arg}_{g,K}.$
For an arbitrary curve $K$, define
${\rm Arg}_{g}(K)=\overline{\lim}_{z\to 0}{\rm Arg}_{g}(z+K).$
Clearly, $|{\rm Arg}_{g}(K)|<\infty$.
@@
Let $f$ be an exponential sum and let $K\subset{\mathbb{C}}$ be a smooth compact curve. Then there exists a constant $C>0$ such that
$|{\rm Arg}_{f}(z+K)|\leq C\qquad\text{for all }z\in{\mathbb{C}}.$
@@
Let $\Lambda$ be the spectrum of $f$, and let $E$ denote the vector space of exponential sums with spectrum $\Lambda$. Denote by $\mathbb{E}$ its projectivization; points of $\mathbb{E}$ correspond to exponential sums defined up to multiplication by a nonzero scalar. Translations by $z\in{\mathbb{C}}$ act on $E$:
$(w\cdot g)(z)=g(z+w),\qquad w\in{\mathbb{C}},$
and hence induce a projective representation
$\rho\colon{\mathbb{C}}\to{\rm Aut}(\mathbb{E}).$
Suppose, to the contrary, that no such constant $C$ exists. Then there is a sequence $\{z_{k}\}\subset{\mathbb{C}}$ such that $|{\rm Arg}_{f}(z_{k}+K)|\to\infty$. Passing to a subsequence, we may assume that $\rho(z_{k})(f)$ converges in $\mathbb{E}$ to a limit $f_{\infty}$. It would then follow that $|{\rm Arg}_{f_{\infty}}(K)|=\infty$, which is impossible. ∎
Proof of Theorem 4.1.
Let $\Delta_{1},\dots,\Delta_{N}$ be the sides of the Newton polygon $\Delta={\rm conv}(\Lambda)$, and let $r_{1},\dots,r_{N}$ be the rays generated by the outward normal vectors to these sides. For each $j$, let $U_{j,R}$ denote the RR–neighborhood of $r_{j}$ in $\mathbb{C}$}. For $R$ sufficiently large, all zeros of $f$ are contained in $\bigcup_{j=1}^{N}U_{j,R}$. It therefore suffices to prove that
$N_{j}(f,r)=\frac{r}{2\pi}\,{\rm length}(\Delta_{j})+O(1),$ (4.2)
where $N_{j}(f,r)$ denotes the number of zeros of $f$ in $U_{j,R}\cap B_{r}$. Let $D_{j}$ be the boundary of the domain $U_{j,R}\cap B_{r}$. By the argument principle,
$N_{j}(f,r)=\frac{1}{2\pi}\,{\rm Arg}_{f}(D_{j}).$
Choose coordinates $x,y$ in $\mathbb{C}$} so that the ray $r_{j}$ coincides with the positive yy–axis. Then $D_{j}$ is the union of four curves: vertical segments $V_{-}(r)$ and $V_{+}(r)$ at $x=\pm R$, an arc $c_{l}$ of the circle of radius $R$ lying below the xx–axis, and an arc $c_{u}(r)$ of the circle of radius $r$ lying above the segment with endpoints $(-R,\sqrt{r^{2}-R^{2}})$ and $(R,\sqrt{r^{2}-R^{2}})$.
For $R$ sufficiently large, the increment of the argument of $f$ along $V_{-}(r)\cup V_{+}(r)$ equals $\sqrt{r^{2}-R^{2}}\,{\rm length}(\Delta_{j})$. The curve $c_{l}$ is independent of $r$, while the increment of the argument along $c_{u}(r)$ is bounded uniformly in $r$ by Lemma 4.1. This proves (4.2) and completes the proof. ∎
@@ 4.2.1. Definitions and examples
(See [21]). A more precise result based on the Fewnomial Theory is given in [22]) Assume that the spectra $\Lambda_{1},\ldots,\Lambda_{n}$ are contained in ${\rm Re\>}({{\mathbb{C}}^{n}}^{*})$. Then, for any domain $D\subset{\rm Im\>}({\mathbb{C}}^{n})$,